put each of the following i y=a(x-h)^2+k form and find the minimum value.
A. f(x)=x^2 +12x +50
B. g(x)=x^2 -6x+3

Answer by pushpaharan(47)

y=a(x-h)^2+k
a- tells you wither you get max or min
if a<0 you will get max
if a>0 you will get min
k- is the value of max or min
h- is the value of the variable wher you will get max or min
(h,k) is the vertex
A)f(x)=x^2 +12x +50
=x^2 +12x + 6^2-6^2+50
=x^2 +12x + 6^2-36+50
=(x+6)^2 + 14
her you get min and value of min =14
B)g(x)=x^2 -6x+3
=x^2 -6x +3^2-3^2 +3
=x^2 -6x +3^2-9 +3
=(x-3)^2 -6
again you will get min with value -6